OpenStudy (anonymous):

On finding the interval of convergence for a power series I need help testing an endpoint. x= -1 for (1/n2^n)*(x-1)^n which would be -2^n/n2^n ....what test would I use? Please show work.

OpenStudy (sirm3d):

\[\large \lim_{n \rightarrow +\infty}\frac{1}{n^{2n}}(-2)^n\]\[\large =\lim_{n \rightarrow +\infty}\frac{(-2)^n}{n^{2n}}\]\[\large =\lim_{n \rightarrow +\infty}(-1)^n\left(\frac{2}{n^{2}}\right)^n\]

OpenStudy (sirm3d):

because this is an alternating series, it is convergent if \[\large \lim_{n \rightarrow +\infty}\left( \frac{ 2 }{ n^2 } \right)^n=0\]divergent otherwise. LET \[\large y=\lim_{n \rightarrow +\infty}\left( \frac{ 2 }{ n^2 } \right)^n\]\[\large \ln y=\lim_{n \rightarrow +\infty} n\ln\left( \frac{ 2 }{ n^2 } \right)\]

OpenStudy (amistre64):

what is the power series?

OpenStudy (anonymous):

Sorry the denominator is not n^(2n), it's n2^n

OpenStudy (anonymous):

power series is \[\sum(1/n2^n)*(x-1)^n\]

OpenStudy (sirm3d):

LOL. so much for the effort. gtg. have work.

OpenStudy (anonymous):

I'm sorry thanks though

OpenStudy (amistre64):

\[\sum \frac{(x-1)^n}{n2^n}\] like this?

OpenStudy (anonymous):

yes

OpenStudy (sirm3d):

\[\large \lim_{n \rightarrow +\infty}\frac{ (-2)^n }{ n 2^n }=\lim_{n \rightarrow +\infty}(-1)^n\frac{ 2^n }{ n 2^n }=\lim_{n \rightarrow +\infty}(-1)^n\frac{ 1 }{ n }=\ln 2\] CONVERGENT!

OpenStudy (anonymous):

but what test is that?

OpenStudy (amistre64):

the steps i take to an interval of convergence is \[\lim_{n\to~inf}\frac{f(n+1)}{f(n)}\] \[\lim_{n\to~inf}\frac{(x-1)^{n+1}}{(n+1)2^{n+1}}~\frac{n2^n}{(x-1)^n}\] \[\lim_{n\to~inf}\frac{n~(x-1)^{n+1-n}}{(n+1)2^{n+1-n}}\] \[\lim_{n\to~inf}\frac{n~(x-1)}{(n+1)2}\] factor out all the no-n parts \[\left|\frac{x-1}{2}\right|~\lim_{n\to~inf}\frac{n}{n+1}\] \[\left|\frac{x-1}{2}\right|*1\] Radiance of convergence is then \[R:\left|\frac{x-1}{2}\right|<1\] \[R:|x-1|<2\] \[R:|x|<3\] giving me an interval of convergence of : -3 < x < 3

OpenStudy (anonymous):

I had an interval of convergence as -1<x<3 which the teacher said is right, the only issue i had was testing the endpoint of x=-1 and what test to use to do that.

OpenStudy (amistre64):

took it a little too far :) R: |x−1| < 2 therefore, -1 < x < 3

OpenStudy (sirm3d):

aternating series test \[\large \sum_{n=1}^{+\infty}(-1)^n a_n\]is convergent if \[\large \lim_{n \rightarrow +\infty}a_n=0\]

OpenStudy (anonymous):

if it was AST then your answer would be divergent because it didn't equal 0....

OpenStudy (amistre64):

well, if you can conform the power series to:\[\ln(\frac{2}{3-x})\] when x=-1 is not a bad construction and the sum when x=-1 in the power series is just 0+0+0+0+... = 0 so it doesnt look like they reach the same value at x=-1 to me; but thats prolly not kosher

OpenStudy (sirm3d):

\[\lim_{n \rightarrow +\infty}\frac{ 1 }{ n }=0\]

OpenStudy (sirm3d):

\[\lim_{n \rightarrow +\infty}(-1)^n \frac{ 1 }{ n }=\ln 2\]

OpenStudy (anonymous):

right so it diverges because it equals ln 2 not 0....but earlier you said that it equaled ln 2 and that converged

OpenStudy (sirm3d):

AST says it is convergent, and the series converges to ln(2)

OpenStudy (anonymous):

no the AST says that the lim of Cn converges if it equals 0, not a number

OpenStudy (sirm3d):

@Autumn0909 you are confused. AST determines only the convergence of an alternating series. The AST tests the SEQUENCE by \[\large \lim_{n \rightarrow +\infty}c_n=0\] and concludes the convergence of the SERIES. It cannot determine the limiting sum of the series.